OpenStudy (anonymous):

-3x^2+3x+2=0 how do you factor it?..... I know the answer becomes (1-x)(3X+2) but im not sure how to get there

OpenStudy (mertsj):

Multiply by -1 to get: 3x^2-3x-2=0 and it is not factorable.

OpenStudy (whpalmer4):

\[(1-x)(3x+2) = 3x+2-3x^2-2x = -3x^2+x+2\]Did you copy part of the problem wrong?

OpenStudy (mertsj):

No he didn't copy it wrong. He forgot to tell you that the right side of the equation is -4x and that the problem has already been solved in another post.

OpenStudy (whpalmer4):

If you had \(-3x^2+x+2=0\) and needed to factor it: First I would eliminate the annoying - sign on the leading term by multiplying through by -1 as suggested. \[3x^2-x-2=0\]Your highest power is \(x^2\) so you'll have two binomials of the form \((ax+b)\) to multiply together to get your polynomial. \[3x^2-x-2 = (ax+b)(cx+d)\] The only factors of 3 are 1 and 3, so we'll take a = 3 and c = 1 \[(3x+b)(x+d) = 3x^2+3dx+bx+bd = 3x^2+(3d+b)x + bd\] Now we need two numbers b,d, that multiply to -2 and if put in (3d+b) = -1 so that our coefficients match. Our only two choices for the multiplication are -1,2 and -2,1. 3(-1)+2=-1, 3(-2)+1=-5, so d=-1 and b = 2. That gives us \[(3x+2)(x+(-1)) = -1(3x+2)(x-1)\] as our factoring, after we restore the factor of -1 we dropped. We could also write it as \[(3x+2)(-x+1) = (3x+2)(1-x)\]if we multiply the (x-1) term by that leading -1.